# Micromegas with GEM preamplification for enhanced energy threshold in low-background gaseous time projection chambers

### v1

@@ Introduction
Expected sensitivity improvement in TREX-DM
As argued in the introduction, there is a strong motivation to extend the sensitivity of dark matter experiments to lower WIMP masses. This requires lowering energy thresholds, as well as reducing the background at the lowest energies.
As of 2022, TREX-DM achieved a background level at low energies of around 80 dru (dru = c keV-1 day-1 kg-1), and an energy threshold around 900 eVee. Reducing the threshold from 900 eVee to 50 eVee significantly enhances sensitivity in the < 1 GeV c-2 region. As already demonstrated in this paper, a great improvement in energy threshold can be achieved by introducing a new electron preamplification stage (the GEM) atop the Micromegas. Laboratory tests (orefsec:results and section 4) have proven GEM extra factors ranging from 20 to 90 are feasible, contingent on gas pressure.
@@ Conclusions
Electron amplification in gas offers an attractive strategy to increase signal to noise ratio and therefore to reduce detector energy thresholds. This feature, coupled with the ability of building radiopure readout Micromegas planes with the microbulk technology, is at the core of the TREX-DM proposal. In this paper we have demonstrated the feasibility of an amplification scheme that should allow to approach the single-electron sensitivity in realistic TREX-DM implementations. The addition of a GEM preamplification stage on top of the microbulk readout, allows for this gain, without jeopardizing the radiopurity specifications of the readout. At the time of writing this article, a GEM + MM combined readout like the one tested in section 4 is being installed and commissioned at the TREX-DM experiment. As discussed in section 5, this improvement would open a new detection window at lower recoil energies that, depending on the background levels achieved at those energies, might lead to substantial improvement to low mass WIMPs, potentially down to unexplored region of the parameter space.
@@
Repository: Data for manuscript submitted to Open Research Europe with title "Micromegas with GEM preamplification for enhanced energy threshold in low-background gaseous time projection chambers"
@@
We acknowledge support from the European Union’s Horizon 2020 Programme, under the European Research Council (ERC), grant agreement ERC-2017-AdG788781 (IAXOplus) and under the Marie Skłodowska-Curie Actions, grant agreement 101026819 (LOBRES). This work is also associated to the IDEAS programme of the 7th EU Framework Programme, grant agreement ERC-2009-StG-240054 (TREX).
We would also like to acknowledge the support from the Agencia Estatal de Investigación (AEI) under the grant agreement PID2019-108122GB-C31 funded by MCIN/AEI/10.13039/501100011033, and under the grant agreement PID2022-137268NB-C51 funded by MCIN/AEI/10.13039/501100011033/FEDER, as well as funds from “European Union NextGenerationEU/PRTR” (Planes complementarios, Programa de Astrofísica y Física de Altas Energías).
Finally, we would also like to acknowledge support from Gobierno de Aragón through their predoctoral research contracts, competitive call 2020-2024.

### v2

@@ Results
Discussion
This section discusses the impact the results presented in section 3 have on the sensitivity of TREX-DM. As argued in the introduction, there is a strong motivation to extend the sensitivity of dark matter experiments to lower WIMP masses. This requires lowering energy thresholds, as well as reducing the background at the lowest energies.
As of 2022, TREX-DM achieved a background level at low energies of around 80 dru (dru = c keV-1 day-1 kg-1), and an energy threshold around 900 eVee. Reducing the threshold from 900 eVee to 50 eVee significantly enhances sensitivity in the < 1 GeV c-2 region. As already demonstrated in this paper, a great improvement in energy threshold can be achieved by introducing a new electron preamplification stage (the GEM) atop the Micromegas. Laboratory tests (subsection 3.1 and subsection 3.2) have proven GEM extra factors ranging from 20 to 90 are feasible, contingent on gas pressure.
@@ Conclusions
Electron amplification in gas offers an attractive strategy to increase signal-to-noise ratio and therefore to reduce detector energy thresholds. This feature, coupled with the ability of building radiopure readout Micromegas planes with the Microbulk technology, is at the core of the TREX-DM proposal. In this paper, we have demonstrated the feasibility of an amplification scheme that should allow to approach the single-electron sensitivity in realistic TREX-DM implementations. The addition of a GEM preamplification stage on top of the Microbulk readout allows for this gain, without jeopardizing the radiopurity specifications of the readout. At the time of writing this article, a GEM + MM combined readout like the one tested in subsection 2.2 is being installed and commissioned at the TREX-DM experiment. As discussed in section 4, this improvement would open a new detection window at lower recoil energies that, depending on the background levels achieved at those energies, might lead to substantial improvement to low mass WIMPs, potentially down to an unexplored region of the parameter space.
@@
Zenodo: Data for manuscript submitted to Open Research Europe with title "Micromegas with GEM preamplification for enhanced energy threshold in low-background gaseous time projection chambers"
@@
This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 788781) (Towards the detection of the axion with the International Axion Observatory (IAXOplus)). This project has also received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No. 101026819 (Low background techniques on particle detectors for Rare Event Searches (LOBRES)). This work is also associated to the IDEAS programme of the 7th EU Framework Programme, grant agreement ERC-2009-StG-240054 (TREX). We would also like to acknowledge the support from the Agencia Estatal de Investigación (AEI) under the grant agreement PID2019-108122GB-C31 funded by MCIN/AEI/10.13039/501100011033, and under the grant agreement PID2022-137268NB-C51 funded by MCIN/AEI/10.13039/501100011033/FEDER, as well as funds from “European Union NextGenerationEU/PRTR” (Planes complementarios, Programa de Astrofísica y Física de Altas Energías). Finally, we would also like to acknowledge support from Gobierno de Aragón through their predoctoral research contracts, competitive call 2020-2024.
@@ Acknowledgments
This article has also been published on arXiv. It is available at https://doi.org/10.48550/arXiv.2412.19864.